### v2

# RefVNLI: Towards Scalable Evaluation of Subject-driven Text-to-image Generation

Abstract
Subject-driven text-to-image (T2I) generation aims to produce images that align with a given textual description, while preserving the visual identity from a referenced subject image. Despite its broad downstream applicability—ranging from enhanced personalization in image generation to consistent character representation in video rendering—progress in this field is limited by the lack of reliable automatic evaluation. Existing methods either assess only one aspect of the task (i.e., textual alignment or subject preservation), misalign with human judgments, or rely on costly API-based evaluation. To address this gap, we introduce RefVNLI, a cost-effective metric that evaluates both textual alignment and subject preservation in a single run. Trained on a large-scale dataset derived from video-reasoning benchmarks and image perturbations, RefVNLI outperforms or statistically matches existing baselines across multiple benchmarks and subject categories (e.g., Animal, Object), achieving up to 6.4-point gains in textual alignment and 5.9-point gains in subject preservation.
[…]
RefVNLI: Towards Scalable Evaluation of
Subject-driven Text-to-image Generation
Aviv Slobodkin1thanks:   Work done during an internship at Google Research.  Hagai Taitelbaum1  Yonatan Bitton1 Brian Gordon1∗Michal Sokolik1Nitzan Bitton Guetta2Almog Gueta1 Royi Rassin1∗Dani Lischinski1Idan Szpektor1 1Google Research   2Ben Gurion University {slobodkin,hagait,yonatanbitton}@google.com
Figure 1: Illustration of RefVNLI: Given a reference image of a subject, a prompt referring to the subject, and a target image, RefVNLI assesses both subject preservation and textual alignment. For subject preservation, it distinguishes identity-preserving variations, like dew on a flower (top image), from identity-altering changes, such as color change (middle image). For textual alignment, it assesses whether the target image reflects all details from the prompt, such as the fence’s position relative to the flower (bottom image).
1 Introduction
In a well-known scene from “The Little Prince”, the narrator attempts to comfort a grieving prince by saying “I’ll draw you a fence around your flower”. While fairly simple, this offer raises a deeper question: what makes such a drawing adequate? Beyond accurately depicting a fence around a flower, the use of ‘your’ implies that it must portray a specific flower—the Prince’s own— one with which he shares a history. Given the flower’s uniqueness and distinct visual traits, the narrator’s task proves far more complex than it first appears.
Figure 2: Qualitative Comparison: We compare RefVNLI with DreamBench++ and CLIP, which score both Subject Preservation (SP) and Textual Alignment (TA), using examples from the Animal, Object, and Human categories. DreamBench++ scores (0-4) are scaled to 0-100 for better readability. RefVNLI exhibits better robustness to identity-agnostic changes (SP), such as the zoomed-out parrot (top-middle) and the zoomed-out person with different attire (bottom-middle). It is also more sensitive to identity-defining traits, penalizing changed facial features (left-most person) and mismatched object patterns (left and middle balloons). Additionally, RefVNLI excels at detecting text-image mismatches (TA), as seen in its penalization of the top-left image for lacking a waterfall.
Subject-driven text-to-image (T2I) generation Chen et al. (2024); Li et al. (2024); Ruiz et al. (2023) enables a variety of downstream applications, such as personalized image generation Ruiz et al. (2023), character consistency in video generation Liu et al. (2024b), and enhancing vanilla T2I evaluation frameworks for less-known entities via image retrieval (Tahmasebi et al., 2025). Unlike standard T2I models that are only conditioned on text inputs, this setup takes both a textual prompt and a reference image, enabling more precise subject representation. For example, when creating an image of a fenced-in flower for the Little Prince, subject-driven models should use a reference image of the Prince’s flower to ensure the output preserves its unique features (see Fig. 1).
Despite its wide applicability, research in this area has been hindered by a lack of reliable and scalable auto-raters. Existing metrics typically correlate poorly with humans, and often focus only on textual alignment between the input prompt and the target image, as in CLIP-T Radford et al. (2021) and SigLIP Zhai et al. (2023), or on subject preservation between the input and target images, as in CLIP-I Radford et al. (2021) and DINO-I Caron et al. (2021), while both aspects are needed for successful subject-driven generation. More correlative metrics, like DreamBench++ Peng et al. (2024) and VIEScore Ku et al. (2024a), depend on costly API calls to models like GPT-4 OpenAI (2024), making them less scalable and reproducible.
To bridge this gap, we present RefVNLI, a cost-effective fine-tuned auto-rater for subject-driven T2I generation. Given a triplet $<$imageref, prompt, imagetgt$>$, RefVNLI predicts two scores—textual alignment and subject preservation—in a single run, as shown in Fig. 1. To train RefVNLI, we automatically curate a large-scale dataset of $<$imageref, prompt, imagetgt$>$ triplets, labeled with $<$textual alignment, subject preservation$>$ $\in\{0,1\}^{2}$. For subject preservation, we identify subjects across video frames, creating positive examples using pairs of frames depicting the same subject, and negative ones by pairing frames of different subjects (yet of the same entity). This approach enables robustness to variations in subject appearance (e.g., rotation, setting, clothing), as well as to the presence of extraneous elements (e.g., dew on the Little Prince’s flower in Fig. 1, top). At the same time, RefVNLI must also be sensitive to identity-defining traits, such as human facial features or object shapes and colors (e.g., middle image in Fig. 1). To this end, we modify images by masking and inpainting identity-critical regions, while keeping everything else unchanged. The original subject crops are then paired with the unaltered images as positive pairs and with the modified images as negative pairs, thereby teaching the model to focus on key identity attributes.
Figure 3: Generating subject preservation classification training instances from video frames. Given two pairs of frames, each extracted from distinct video scenes featuring the same entity (e.g., a dog), where both frames within each pair depict the same subject (e.g., the same dog), we construct training {imageref, imagetgt} pairs for subject preservation classification. Positive pairs are formed by pairing a cropped subject from one frame (e.g., dog from left frame in Scene 1) with the full frame from the same scene (right frame in Scene 1). In contrast, negative pairs are created by pairing the cropped subject with the other scene’s full frames (e.g., Scene 2). This process is applied to all four frames, with each taking turns as the cropped reference image (imageref), while the corresponding full-frame counterparts serve as imagetgt, yielding a total of 4 positive and 8 negative training pairs.
For textual alignment, we first create positive image-prompt pairs. For that, we use an LLM to caption each image in the aforementioned pairs, ensuring focus on the subject by enclosing it within a bounding box. Negative pairs are then formed by replacing these captions with those of different scenes. For extra sensitivity to minor mismatches, like a fence drawn next to rather than around the Prince’s flower (Fig. 1, bottom), we also create negative pairs by altering a single fact in each original (positive) caption. Finally, to derive the <<<imageref, prompt, imagetgt>>> triplets from each pair of frames and associated captions, we use the cropped subject from one frame as imageref and the entire second frame, alongside its caption, as {prompt, imagetgt}, resulting in a total of 1.2 million instances.
We evaluate RefVNLI on multiple human-labeled test sets for subject-driven generation, including DreamBench++ Peng et al. (2024), ImagenHub Ku et al. (2024b), and KITTEN Huang et al. (2024), across categories such as Humans, Animals, Objects, Landmarks, and a multi-subject setting. For textual alignment, RefVNLI consistently matches or outperforms all baselines, with up to 6.4-point gains in Landmarks and proficiency at detecting subtle text-image misalignments (e.g., missing waterfall in Fig. 2, top-left). It also leads in subject preservation, with gains of up to 6.3 points on Objects (Table 1) and 5.9 points in the multi-subject setting, surpassing the larger GPT-4o-based DreamBench++ baseline. As seen in Fig. 2, it balances robustness to non-critical changes (e.g., zoomed-out toucan, top-middle) with sensitivity to identity shifts (e.g., altered facial features, bottom-left). Further, RefVNLI effectively handles rare subjects (§5), outperforming all baselines and highlighting its value as a reliable alternative to standard T2I metrics for uncommon entities.
2 RefVNLI: Automatic Metric for Subject-driven T2I Generation
We introduce RefVNLI, a cost-effective auto-rater specifically tailored for subject-driven T2I generation. This section details the automated pipeline used to construct its training dataset (§2.1) and the subsequent training process of RefVNLI2.2).
2.1 Training Dataset Construction
To train RefVNLI, we collect a large scale dataset of <<<imageref, prompt, imagetgt>>> triplets, each with two binary labels: one for subject preservation of imageref in imagetgt, and one for textual alignment between the prompt and imagetgt. This involves first creating subject-driven {imageref, imagetgt} pairs, followed by automatic generation of subject-focused prompts for each imagetgt.
Figure 4: Creating identity-sensitive {imageref, imagetgt} pairs. Starting with an image and a mask of a subject (e.g., a bag), we randomly keep 5 patches within the masked area ([1]) and use them to create 5 inpainted versions ([2]). The version with the highest MSE between the altered and original areas (e.g., bottom image, MSE = 3983) is paired with the unmodified crop to form a negative pair, while the original image and the same crop create a positive pair, with the crop acting as imageref in both cases.
Subject-driven image pairs.
To ensure our {imageref, imagetgt} dataset is robust to identity-agnostic changes (e.g., pose, clothing, or lighting changes), we use video-based datasets that inherently capture these differences. Specifically, we use Mementos Wang et al. (2024), comprising scene-specific video frames with human-written textual descriptions, and TVQA+ Lei et al. (2020), containing human-annotated bounding boxes for characters and objects in TV episodes. We first locate subjects within frames: for Mementos, we extract entities from the provided textual descriptions using Gemini Team (2024) and localize them in the associated frames with an object detection model Minderer et al. (2022), while for TVQA+, we directly use the provided bounding boxes. Positive pairs are formed from frames featuring the same subject, usually within the same scene, [^1] For TVQA+, we also include cross-scene positive pairs for named entities, such as TV characters. while negative pairs consist of frames with distinct subjects (of the same type of entity), often across scenes (see Fig. 3). These frame-pairs are then converted into {imageref, imagetgt}-pairs by cropping the subject from one frame as imageref (e.g., left frame in Fig. 3, Scene 1) and using the full second frame as imagetgt (each of the other frames in Fig. 3). This is then repeated with reversed roles for an extra {imageref, imagetgt} pair. In total, we collected 338,551 image pairs (228,661 from Mementos and 109,890 from TVQA+) from 44,418 unique frames.
To further enhance sensitivity to identity-specific attributes, such as facial features in humans or shapes and patterns in objects, we leverage the Open Images dataset Kuznetsova et al. (2020) to create additional training instances, as shown in Fig. 4. Using its gold segmentation masks, we selectively mask and inpaint identity-critical regions while preserving other details. Specifically, we randomly sample 5 sub-masks covering 30%-50% of the subject mask ([1] in Fig. 4), which we use to create 5 inpainted variants ([2]). The version with the highest Mean Squared Error (MSE) between the modified and original regions (e.g., Fig. 4, bottom image, MSE=3983) is then paired with the unmodified cropped subject to form a negative pair of {imageref, imagetgt}, while the original image and the same crop form a positive pair, with the crop serving as imageref in both cases. This process yields extra 16,572 pairs, helping the model focus on fine-grained identity details. To further improve data quality, we also apply multiple filtering steps, including removing blurry images and those with unclear subjects (see Appendix C.2 for details).
Image-prompt pairs.
For each {imageref, imagetgt} pair, we generate positive and negative prompts for imagetgt (Fig. 5). Positive prompts (Fig. 5, top) are created by instructing Gemini Team (2024) to describe imagetgt, ensuring the subject is explicitly mentioned by enclosing it in a bounding box and guiding the model to focus on it, as well as filtering out prompts lacking it. For negative prompts (Fig. 5, middle), we swap prompts between frames containing the same entity type (e.g., a dog). To further enhance sensitivity to subtle mismatches, we also create hard-negative prompts (Fig. 5, bottom) by using Gemini to modify a single non-subject detail in the positive prompts, following Gordon et al. (2024). In total, this and the image-pairing steps yield 1.2 million <<<imageref, prompt, imagetgt>>> triplets labeled for textual alignment and subject preservation.
2.2 RefVNLI Training
[…]
Figure 5: Example of prompt-imagetgt pairs. Given an image with some subject (e.g., a dog), we create a positive prompt by adding a bounding box around the subject and directing Gemini to describe it (top prompts). Negative prompts are created by swapping prompts between images of the same entity (middle prompts). For additional hard negatives, we guide Gemini to modify a single non-subject detail in the positive prompt while keeping the rest unchanged (bottom prompts).
3 Experimental Settings
This section outlines our meta-evaluation protocol and benchmarks (§3.1), followed by an overview of the baseline models used for comparison (§3.2).
3.1 Meta-evaluation and Benchmarks
We include 3 subject-driven generation benchmarks with human annotations for textual alignment and subject preservation across categories such as Human, Animal, Object, and Landmark. To enable a unified evaluation framework, given differing scoring methods (5-scale and binary), we convert all annotations into binary labels: one for whether imagetgt fully captures the prompt (textual alignment) and another for whether it correctly depicts the referenced subject (subject preservation).
For meta-evaluation, we report ROC AUC for each criterion, following standard practice Honovich et al. (2022); Yarom et al. (2023); Zha et al. (2023), and also compute a unified score as the harmonic mean of the two scores. Following Honovich et al. (2022), significance testing is assessed via bootstrap resampling (Efron, 1987), comparing each baseline to RefVNLI. We report mean scores and highlight models with statistically significant under- or outperformance relative to RefVNLI.
We next present the 3 analyzed benchmarks.
Dreambench++
Peng et al. (2024) is a subject-driven generation benchmark with human annotations for 8,190 images generated by 7 models. Annotators rated textual alignment and subject preservation on a 0-4 scale, with each image evaluated by 2 raters. To convert these ratings into binary labels, we classify a criterion as positive if both scores are at least 3, and at least one is a 4. We report performance separately for the benchmark’s three subject categories: Human, Animal, and Object. [^4] A fourth ‘style’ category is excluded as it is beyond our work’s scope.
Textual Alignment Subject Preservation Unified Evaluation
Animal Human Object Animal Human Object Animal Human Object
CLIP   72.8   77.4   74.6   72.4 87.7   76.4   72.6 82.2   75.5
DINO - - - 80.1   78.0   77.3 - - -
Crop-IR - - -   76.9 85.6 83.4 - - -
ArcFace - - - -   61.0 - - - -
CLIPScore   71.5   76.1   72.9 - - - - - -
BLIPScore   75.4   79.5   78.9 - - - - - -
SigLIP   72.5 80.2   77.1 - - - - - -
TIFA   70.6   75.7   69.5 - - - - - -
VQAScore 79.4   78.0 82.6 - - - - - -
VIEScore 77.9   75.2   73.3   63.4 81.1   76.4   69.8 77.7   74.8
DreamBench++ 79.5 82.7 82.5   74.5 84.1   79.4   76.9 83.4   80.9
[…]
RefVNLI 80.2 82.5 82.0 79.4 86.0 85.7 79.8 84.2 83.8
Table 1: ROC AUC scores on DreamBench++ for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation) across Animal, Human, and Object categories. The last two rows feature models finetuned on our dataset, with PaliGemmatext/ref comprising two separate models (PaliGemmatext and PaliGemmaref) trained exclusively for each criterion. Bold indicates the highest score per column. and indicate statistically significant underperformance and outperformance relative to RefVNLI, respectively.
ImagenHub
(Ku et al., 2024b) is a human-annotated benchmark for conditional image generation, covering a subject-driven task (150 instances) and a multi-concept task (102 instances), which involves 2 referenced subjects per instance. Each image was rated by 3 annotators. Instead of separate ratings for textual alignment and subject preservation, annotators provided a single adherence score (0, 0.5, or 1) per image. To align with our binary labeling framework, images rated 1 by all 3 annotators were assigned positive labels for both criteria, while the rest were re-annotated by this paper’s authors. In the Multi-subject setting, a positive subject preservation label was assigned only when both subjects were accurately depicted. For evaluation, we report separate scores for Animals and Objects in the single-subject task, while Multi-subject instances are split into two single-subject evaluations, with the final score being the lower rating (per criterion), to ensure a stricter assessment.
KITTEN
Huang et al. (2024) evaluates subject-driven T2I models on generating diverse real-world entities (e.g., plants, vehicles, landmarks), using 5 reference images and a prompt. Annotators rated entity depiction on a 1–5 scale and provided binary textual alignment scores, with each image assessed by 5 annotators. Unlike our focus on specific subjects, KITTEN evaluates general entity alignment (e.g., a generic rose rather than a specific one). Hence, we only use the 256 Landmark images, as landmarks are unique entities where entity adherence coincides with subject adherence. To convert ratings into binary labels, we apply majority voting for textual alignment and consider subject preservation positive only if most annotators rated it at least 4 and the average score is 4 or higher.
3.2 Baselines
We evaluate RefVNLI against both standard and state-of-the-art methods for measuring textual alignment, subject preservation, or both.
Baselines for textual alignment.
We compare RefVNLI with two groups of automatic metrics for textual alignment. The first group leverages large vision-language models (VLMs), computing cosine similarity between text and image encodings. This includes BLIPScore Li et al. (2022), CLIPScore Hessel et al. (2021), and SigLIP Zhai et al. (2023). The second group, which includes TIFA Hu et al. (2023) and VQAScore Lin et al. (2024), evaluates textual alignment via visual question answering (VQA). We also include a baseline where PaliGemma is finetuned on our dataset exclusively for textual alignment, given only the prompt and target image, referred to as PaliGemmatext.
Baselines for subject preservation.
For subject preservation, we compare RefVNLI to baselines that use large VLMs by computing cosine similarity between reference and target image embeddings. These include DINO Caron et al. (2021), Crop-IR Winter et al. (2024), [^5] See Appendix E for more details. and for the Human category, also ArcFace Deng et al. (2019), a face-recognition model. We also assess a PaliGemma model finetuned on our dataset solely for subject preservation, using only reference and target images (formatted as in §2.2), denoted as PaliGemmaref.
Baselines for both criteria.
We also include 3 metrics that assess both criteria. CLIP Radford et al. (2021) computes scores separately for each criterion by calculating cosine similarity between the encodings of imagetgt and those of prompt and imageref. VIEScore Ku et al. (2024a) uses an elaborate GPT-4o OpenAI (2024) few-shot strategy, simultaneously generating two 0–10 ratings, one for each criterion. Lastly, DreamBench++ Peng et al. (2024) evaluates each criterion separately using distinct GPT-4o prompts with hand-crafted instructions and examples. This method follows a two-step prompting process, where GPT-4o first summarizes the evaluation task to increase task comprehension before assigning a 0–4 score.
Textual Alignment Subject Preservation Unified Evaluation
Animal Object Multi-subj. Animal Object Multi-subj. Animal Object Multi-subj.
CLIP 81.8   74.7   81.1   63.8   73.3   52.6   71.6   74.0   63.8
DINO - - - 81.7   77.3   50.0 - - -
Crop-IR - - - 77.6 84.1 56.8 - - -
CLIPScore 81.5   75.0   79.1 - - - - - -
BLIPScore 82.9   79.7   84.2 - - - - - -
SigLIP 80.7   80.6   82.3 - - - - - -
TIFA 79.9   76.1   79.2 - - - - - -
VQAScore   77.3   83.8 87.8 - - - - - -
VIEScore   62.1   54.1   71.6   56.4   49.4   50.2   59.0   51.5   58.9
DreamBench++ 86.4   85.5 88.2   71.1 84.0   54.3   78.0 84.8   67.2
[…]
RefVNLI 84.6 89.4 86.2 80.2 83.8 62.7 82.3 86.5 72.6
Table 2: ROC AUC scores on ImagenHub for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation) across Animal and Object categories, as well as for the Multi-subject setting.
Textual Alignment Subject Preservation Unified Evaluation
CLIP   83.2 80.1   81.5
DINO - 85.4 -
Crop-IR - 90.2 -
CLIPScore   83.3 - -
BLIPScore   82.6 - -
SigLIP   75.3 - -
TIFA   90.6 - -
VQAScore   89.0 - -
VIEScore   82.5 87.5 84.9
DreamBench++   87.0   89.9 88.4
[…]
RefVNLI 97.0 82.2 88.9
Table 3: ROC AUC scores on KITTEN (landmarks) for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation).
4 Results
Our main results are summarized in Tables 1, 2, and 3, with qualitative examples in Fig. 2.
On DreamBench++ (Table 1), RefVNLI outperforms or statistically matches all baselines across both criteria, with a notable 6.3-point lead over the GPT-4o-based DreamBench++ metric in subject preservation for Objects. This is especially notable given the benchmark’s diverse visual styles, including cartoonish and pixelated images, which are outside RefVNLI’s training distribution of real-world video frames. Similarly, on ImagenHub (Table 2), RefVNLI matches or exceeds all baselines in both single- and multi-subject settings, with 5.9-point gains over the strongest non-finetuned model on subject preservation of the multi-subject setting (Crop-IR). Lastly, on KITTEN (Table 3), RefVNLI leads in textual alignment but underperforms in subject preservation, though it remains statistically comparable to most baselines. This may result from RefVNLI’s identity-sensitive training, which penalizes minor deviations—especially challenging for landmarks with intricate visual details (Fig. 12)—and from a domain shift, as landmarks were absent from RefVNLI’s training data (OOD).
Notably, across all benchmarks, fine-tuning only for textual alignment (PaliGemmatext) slightly reduces performance, especially for Animals and Humans, while training solely for subject preservation (PaliGemmaref) yields even larger declines—up to a 14.8 points for Humans (Table 1). This suggests that joint training provides complementary benefits, with subject preservation gaining the most.
Fig. 2 further showcases RefVNLI’s strengths, like its sensitivity to subtle textual alignment errors, such as a missing waterfall (top-left). For subject preservation, it remains robust to identity-agnostic changes, like a zoomed-out parrot or person (top-center and bottom-center) or different clothes (bottom-center), while staying sensitive to key identity traits, e.g., changed facial features (bottom-left) and colors (left and middle balloons).
Overall, RefVNLI consistently outperforms or statistically matches all baselines on both criteria, with the only exception of subject preservation in the OOD landmarks category, where it still performs competitively. Importantly, it offers the best trade-off between textual alignment and subject preservation, surpassing all non-finetuned metrics in Unified Evaluation across all benchmarks.
5 Applicability to Rare Entities
To test RefVNLI on unfamiliar subjects, we use the ImageRAG benchmark Shalev-Arkushin et al. (2025), which evaluates generated images based on prompts and reference images of uncommon subjects (e.g., scientific animal names, lesser-known dishes). Human annotators compared image pairs, selecting the better one based on Textual Alignment, Visual Quality (evaluating general depiction of the entity rather than exact reference-adherence), and Overall Preference. We report per-axis accuracy, defined as the frequency with which a metric ranks the human-preferred image higher, with approximate ties handled by rounding. Overall Preference is computed as the harmonic mean of textual and visual scores, and significance testing is assessed via bootstrap resampling, akin to §3.1.
As shown in Table 4, RefVNLI consistently outperforms all baselines in aligning with human preferences across these criteria, showcasing strong robustness to rare subjects. [^6] TIFA was excluded due to assigning identical scores to 61% of pairs, making accuracy calculations unreliable. This is further supported by Fig. 6, where only RefVNLI repeatedly matches human selections.
Textual Alignment Visual Quality Overall Preference
CLIP   51.8 91.3   69.2
DINO - 91.4 -
Crop-IR -   86.4 -
CLIPScore   47.4 - -
BLIPScore   39.6 - -
SigLIP   74.8 - -
VQAScore   52.3 - -
VIEScore   60.8   65.4   69.6
DreamBench++   56.6   83.1   78.9
[…]
RefVNLI 87.2 95.5 91.4
Table 4: Results on ImageRAG rare concepts, where users select the better image in each pair based on textual alignment, visual quality (general entity depiction rather than specific subject adherence), and overall preference. We report accuracy: how often models ranked the human-preferred image higher. Overall preference is the harmonic mean of textual and visual scores.
6 Related Work
Evaluation of Visual Language Models (VLMs) spans various settings, including visual reasoning Bitton-Guetta et al. (2024); Kahou et al. (2017) and visual question-answering Antol et al. (2015); Marino et al. (2019); Mensink et al. (2023). For text-to-image (T2I) models, assessments normally focus on image quality Heusel et al. (2017); Salimans et al. (2016), diversity Rassin et al. (2024), and alignment with the text Hessel et al. (2021); Radford et al. (2021); Hu et al. (2023); Yarom et al. (2023); Zhai et al. (2023); Lin et al. (2024). Assessing subject preservation, which is crucial for subject-driven generation, is typically done using embedding-based metrics like CLIP Radford et al. (2021) and DINO Caron et al. (2021). Other metrics, like VIEScore Ku et al. (2024a) and DreamBench++ Peng et al. (2024), use GPT-4o OpenAI (2024) to measure both criteria.
Subject-driven T2I models have been gaining much traction, with some methods fine-tuning general models into specialist versions that capture specific subjects and styles Gal et al. (2022); Kumari et al. (2023); Ruiz et al. (2023); Sohn et al. (2023); Park et al. (2024). Others focus on broader applicability using one-shot examples, either through adapter-based methods that integrate encoded reference images into diffusion models Gal et al. (2023); Jia et al. (2023); Wei et al. (2023); Ye et al. (2023) or via adapter-free techniques that directly use extracted features such as attention maps Liu et al. (2023); Hertz et al. (2024); Lv et al. (2024).
Closely related, image editing complements subject-driven T2I generation in that the generated image’s appearance is primarily governed by the input image, with the text only impacting specific aspects, whereas in out setting it is the other way around. The task has evolved from pixel-to-pixel translation for predefined transformations Isola et al. (2017); Zhu et al. (2017); Wang et al. (2018) to more flexible, text-guided edits Brooks et al. (2023); Tumanyan et al. (2023); Parmar et al. (2023), with recent diffusion-based methods improving precision via cross-attention manipulation Hertz et al. (2022); Yang et al. (2023). Beyond images, personalized generation extends to other modalities, including videos and texts. Video generation can be conditioned on text Li et al. (2018); Hong et al. (2022); Singer et al. (2022), reference images Wei et al. (2024); Zhou et al. (2024), or other videos Ku et al. (2024c). In text generation, efforts focus on style transfer Reif et al. (2022); Zhang et al. (2024), debiasing Zhao et al. (2018); Ravfogel et al. (2020), and broader semantic control Shapira et al. (2022); Slobodkin et al. (2023); Xie et al. (2023).
Figure 6: ImageRAG Rare Entities Examples: We compare RefVNLI with CLIP and DreamBench++ in aligning with human preferences (top rows of each example) across Textual Alignment (TA), Image Quality (IQ), and Overall Preference (OP). DreamBench++ scores (0–4) are rescaled to 0–100 for readability. The higher of the two criterion-wise scores is emphasized unless both are equal.
Finally, several studies leveraged intra-frame relationships in videos to learn more human-aligned visual representations. These works aim to improve robustness to identity-agnostic variations (e.g., rotation, lighting), by analyzing consecutive frames sourced from public video datasets Jin et al. (2018); Parthasarathy et al. (2023); Wang and Gupta (2015); Wang et al. (2017); Wu and Wang (2021) or captured by cameras on moving agents Agrawal et al. (2015); Jayaraman and Grauman (2015).
7 Conclusion
We present RefVNLI, a cost-effective and reliable metric for subject-driven T2I evaluation that jointly assesses textual alignment and subject preservation. Trained on a large-scale, auto-generated dataset, RefVNLI is designed to be robust to identity-agnostic visual variations (e.g., pose, lighting, background) while remaining sensitive to identity-specific features (e.g., facial features, object shape, and unique details) when evaluating subject preservation. For textual alignment, it leverages subject-specific prompts with perturbed hard negatives to detect and penalize fine-grained mismatches. Across benchmarks, RefVNLI outperforms or rivals all baselines, including larger GPT-4o-based metrics, particularly on less-common subjects.
Future work should focus on improving performance across artistic styles as well as identity-altering edits, and supporting multiple reference images. More broadly, RefVNLI facilitates progress in personalized T2I generation by enabling better checkpoint selection, reinforcement learning, and iterative model refinement.
8 Limitations
RefVNLI was trained on data sourced from real-life video frames and images. While it performs well on stylistically consistent inputs, including cartoonish or pixelated images, it struggles with cross-style scenarios where imageref and imagetgt differ in style, as well as when the subject undergoes explicit modifications (e.g., changes in color or shape). Additionally, the current framework is limited to single-reference cases and should be extended to support multiple references, both for the same subject and for distinct ones.
Moreover, research on subject-driven generation could benefit from a unified score capturing overall performance, rather than the two separate scores currently provided by RefVNLI. Although the harmonic mean of the two offers a reasonable proxy, future iterations should aim to output a single, integrated metric, alongside the individual, more granular scores.
Appendix A Reproducibility
[…]
At inference, we used the same input structure, including the insertion of markups around the subject in the prompt and separate image inputs. The textual alignment and subject preservation scores were then computed from the probabilities of the first and second generated tokens, respectively, with each score defined as the probability of token ‘1’ divided by the sum of probabilities for tokens ‘0’ and ‘1’.
[…]
Appendix B Additional Details on the Meta-evaluation
For meta-evaluation, we perform bootstrapping (Efron, 1987) (1,000 samples per benchmark, with repetitions). For the ImageRAG rare entities benchmark, consisting of only 26 instances, we sample each time a sample 4 times this size, to ensure there are at least 100 instances. For the other benchmarks, which are significantly larger, the size of each sample is identical to the original size of the corresponding benchmark. For the calculations of confidence intervals (CIs), we use significance level $p=0.05$.
Context
Misalignment Injection Instructions (Short Captions) 1. Understand the Caption: Carefully read the short caption to fully grasp the scene it describes. 2. Identify and Swap: Select a single visual detail within the caption to modify. Replace this detail with a different, incorrect, but still plausible visual detail. For example, you might change a color, an object, or a location. Do not modify the underlined entity (if any). 3. Apply the Tags: Enclose the original visual detail within <swap> tags. Immediately after the closing </swap> tag, write the new, incorrect visual detail. There should be no space between the closing </swap> and the new word. Example: If the original sentence is "The cat sat on the red mat," and you want to change "red" to "blue," the result should be: "The cat sat on the <swap>red</swap><blue> mat." 4. Final Check: Ensure the modified caption is grammatically correct and reads naturally, even though it now contains a factual error. The sentence should be internally logical, despite contradicting the actual visual content. Again, ensure the underlined entity (if any) remains completely unchanged.
Few-Shot
Here are some examples:
INPUT: A woman is sitting in a living room, and <u>she</u> is looking at something with a concerned expression
OUTPUT: A woman is sitting in a </swap>living room</swap><kitchen>, and <u>she</u> is looking at something with a concerned expression.
INPUT: Two men are sitting on a leather couch in a living room. One <u>man</u> is sitting on the left side of the couch, looking at a laptop. The other man is sitting on the right side of the couch, talking on a phone. The room is decorated with various items, including a large model of a spaceship.
OUTPUT: Two men are sitting on a leather couch in a living room. One <u>man</u> is sitting on the left side of the couch, looking at a laptop. The other man is sitting on the right side of the couch, talking on a phone. The room is decorated with various items, including a large model of a <swap>spaceship</swap><sailboat>.
Now it’s your turn! Follow the instructions. Answer only with the corrupted sentence, Don’t forget to add the tags.
INPUT: A lizard is perched on a rock, surrounded by other rocks and foliage. The <u>lizard</u> is facing the camera, with its head raised and its tail curled behind it.
OUTPUT:
Generated
OUTPUT: A lizard is perched on a <swap>rock</swap><branch>, surrounded by other rocks and foliage. The <u>lizard</u> is facing the camera, with its head raised and its tail curled behind it.
Figure 8: Hard Negative Caption Generation. This figure illustrates the prompting strategy used to generate hard negative captions, containing a single, plausible but factually incorrect visual detail, for enhanced misalignment detection.
Appendix C Further Information on the Data Construction Pipeline
C.1 Collection of Subject-driven Image Pairs
To reduce noise when collecting subject-driven image pairs, we applied several filtering steps, including the removal of blurred images and those not depicting the intended subject. Subject presence was verified using Gemini (Team, 2024) (version gemini-2.0-flash). For the subset sourced from TVQA+, we additionally filtered out frames containing subtitles or credits, also using Gemini.
For identity-sensitive image pairs, we used Stable Diffusion (Rombach et al., 2022) for inpainting, [^7] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting with $\eta=1.0$ and a guidance scale of 3.0. We retained only images with a full mask size of at least 60,000 pixels (20,000 for humans, focusing on facial regions). Five patches of 250–300 pixels were randomly sampled and inpainted. To increase the likelihood of meaningful subject changes, we further filtered out inpainted images where all patchwise MSE values fell below 6,500 for objects, 5,400 for animals, and 20,000 for humans.
C.2 Collection of Image-prompt Pairs
For the image-captioning of the imagetgt with the inserted bounding boxes, we employed Gemini (Team, 2024) (version gemini-2.0-flash).
C.3 Generation of prompt-imagetgt Hard-negatives
Fig. 8 showcases the prompt used to generated the prompt-imagetgt hard negatives.
Textual Alignment Subject Preservation Unified Evaluation
DreamBench++ ImagenHub (Single/Multi) KITTEN DreamBench++ ImagenHub (Single/Multi) KITTEN DreamBench++ ImagenHub (Single/Multi) KITTEN
RefVNLI (ours) 81.5 87.7 / 86.3 97.0 82.7 83.0 / 62.8 82.3 82.1 85.3 / 72.7 89.0
      reverse classification order 80.0 85.2 / 85.5 95.3 80.9 84.3 / 68.7 87.0 80.4 84.7 / 76.2 91.0
      multiclass 79.5 83.7 / 84.7 94.7 79.6 76.0 / 61.1 86.3 79.5 79.7 / 71.0 90.3
      separate classification 79.7 85.2 / 87.5 95.8 78.3 77.1 / 56.7 89.2 79.0 80.9 / 68.8 92.4
      no markup 78.4 87.0 / 84.3 92.3 65.5 75.9 / 60.8 88.7 71.4 81.1 / 70.6 90.5
      concatenated images 79.6 86.2 / 86.2 93.6 74.2 81.1 / 81.2 89.9 76.8 83.6 / 83.6 91.7
      only video-based training 80.4 85.2 / 83.8 96.3 77.6 77.5 / 66.4 89.7 79.0 81.2 / 74.1 92.9
Table 5: Ablation Study: ROC AUC scores for various ablated versions of RefVNLI across benchmarks (over all subjects). The ablations evaluate alternative output formulations, such as reversed classification order, a four-label multiclass framework, and separate aspect classification via a special token. Additional variants exclude subject markup from input prompt, merge reference and target images, or remove identity-sensitive training examples by using only video-based instances.
Appendix D Ablations
To assess the impact of various design decisions in RefVNLI, we run an ablation study examining alternative input and output configurations. On the output side, we test: reversing the classification order (subject preservation before textual alignment); a 4-label multiclass framework for joint text-image alignment classification; and a model that prefixes a designated token (‘TEXT’ or ‘IMAGE’) to the prompt to enable separate classification of each aspect within a unified model. For inputs, we explore the effect of removing subject markup from the prompt and of concatenating imageref and imagetgt instead of passing them separately. Finally, we also explore the impact of omitting the identity-sensitive training examples (see §2.1), by only using the video-based instances during training.
Results (Table 5) show that reversing the classification order degrades performance, particularly in subject preservation, as does evaluating each aspect separately. This suggests that first evaluating textual alignment helps in subject preservation assessment. The multiclass approach also underperforms compared to our dual binary classification setup, highlighting the benefits of treating each criterion independently. Further, removing subject markup weakens subject preservation, underscoring its role in linking the reference image to the prompt. Additionally, concatenating images instead of processing them separately harms performance, emphasizing the advantage of distinct image inputs. Finally, excluding identity-sensitive training instances leads to notable drops in subject preservation, underscoring their importance.
DreamBench++ ImagenHub KITTEN
Animal Human Object Animal Object Multi-subj. Landmarks
OWL-ViT 77.4 87.1 81.8 74.8 83.1 52.0 85.6
GroundingDINO 76.9 85.6 83.4 77.6 84.1 56.8 90.2
Table 6: ROC AUC scores of Crop-IR for subject preservation when employed with two different object-detection models: OWL-ViT and GroundingDino. Bold indicates the highest score per column.
Appendix E Crop-IR Object Detection Model Ablation
Deploying the Crop-IR metric (Winter et al., 2024) requires an object detection model to locate and crop the referenced subjects. To this end, we compare two prominent object detection models: OWL-ViT (Minderer et al., 2022) [^8] https://huggingface.co/google/owlv2-base-patch16 and GroundingDino (Liu et al., 2024a). [^9] https://huggingface.co/IDEA-Research/grounding-dino-base As shown in Table 6, GroundingDino leads to better evaluation of subject preservation, and is therefore adopted to ensure a fairer evaluation.
Inference Time (seconds) GPU Memory Usage (GiB) API Calls Cost (\$)
CLIP 0.1 1.2 -
DINO 0.06 0.7 -
Crop-IR 0.6 5.8 -
ArcFace 1.2 - -
CLIPScore 0.07 0.6 -
BLIPScore 0.7 4.5 -
SigLIP 0.03 1.4 -
TIFA 22.5 26.4 -
VQAScore 0.2 23.1 -
VIEScore 6.9 - 0.04
DreamBench++ (text) 3.2 - 0.02
DreamBench++ (ref) 1.0 - 0.03
[…]
PaliGemmaref 0.4 12.5 -
RefVNLI 0.5 12.5 -
Table 7: Computational costs for all baseline models and RefVNLI, including per-instance inference time (in seconds), GPU memory usage (in GiB), and GPT-4o API costs (in \$, only when applicable), averaged across benchmarks. For DreamBench++, we report separate values for each evaluation criterion, as each requires a distinct API call under its framework. The final three rows present models fine-tuned on our dataset, with PaliGemmatext and PaliGemmaref being the variants tuned exclusively for evaluating textual alignment and subject preservation, respectively.
Appendix F Computational Cost
Table 7 presents the computational costs of all baseline models and RefVNLI, in terms of inference time, GPU memory usage and GPT-4o API costs.
Appendix G Additional Qualitative Examples for Subject Preservation Evaluation
In Figures 9, 10, 11, and 12 we present additional qualitative examples of subject preservation evaluation for the Animal, Human, Object, and Landmark categories, respectively.
Figure 9: Qualitative Examples of Subject Preservation Evaluation for the Animal Category. DreamBench++ scores (0-4) are scaled to 0-100 for better readability.
Figure 10: Qualitative Examples of Subject Preservation Evaluation for the Human Category. DreamBench++ scores (0-4) are scaled to 0-100 for better readability.
Figure 11: Qualitative Examples of Subject Preservation Evaluation for the Object Category. DreamBench++ scores (0-4) are scaled to 0-100 for better readability.
Figure 12: Qualitative Examples of Subject Preservation Evaluation for the Landmark Category. DreamBench++ scores (0-4) are scaled to 0-100 for better readability.
Appendix H List of Data and Software Licenses Employed in this Paper
Our framework dependencies are:
  1. 1.
    Mementos dataset: https://github.com/si0wang/Mementos, Misc.
  2. 2.
    TVQA+ dataset: https://github.com/jayleicn/TVQAplus/blob/master/LICENSE, under the MIT License.
  3. 3.
    Open Images dataset: https://github.com/openimages/dataset/blob/main/LICENSE, under an Apache License 2.0.
  4. 4.
    PaliGemma model: https://ai.google.dev/gemma/terms, under Gemma Terms of Use License.
  5. 5.
    Gemini model: https://ai.google.dev/gemini-api/docs/models, under an Apache License 2.0.
  6. 6.
    GPT-4o model: https://github.com/openai/openai-openapi/blob/master/LICENSE, under the MIT License.
  7. 7.
    OWL-ViT model: https://huggingface.co/google/owlv2-base-patch16, under an Apache License 2.0.
  8. 8.
    GroundingDino model: https://huggingface.co/IDEA-Research/grounding-dino-base, under an Apache License 2.0.
  9. 9.
    Stable Diffusion inpainting model: https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0/blob/main/LICENSE.md, under Stability AI CreativeML Open RAIL++-M License.